OpenStudy (anonymous):

The diameter of the larger circle is 20.4 cm. The diameter of the smaller circle is 10.8 cm. What is the approximate area of the shaded region? A. 326.7 cm² B. 289.4 cm² C. 235.1 cm² D. 72.3 cm²s

OpenStudy (anonymous):

find the area of the larger circle. find the area of the smaller circle... then, subtract the two areas..

OpenStudy (harsha19111999):

Area of circle = \[\pi r^2\] So subtract the areas

OpenStudy (anonymous):

note: the given are the diameters so find first the radius :)

OpenStudy (anonymous):

The radius for the shades circle would be 10.2 The radius for the unshaded circle would be 5.4

OpenStudy (anonymous):

subtract the larger circles area from smalller one`s . u will get the answer

OpenStudy (anonymous):

area of larger circle - area of smaller circle = \[\pi(10.2)^2 - \pi(5.4)^2\] = \[104.04\pi - 29.16\pi \] = \[78.88\pi \] = 235.1 ...